OpenStudy (anonymous):

Please help teach me to solve this question on parallel vectors (Thankyou in advance): Consider the points P(-3,1,0,5), Q(0,2,1,-3), and R(-2,1,4,0). Find the point S in R4 whose third component is 8 and such that PQ is parallel to RS. S = ( ?, ?, 8, ?)

OpenStudy (asnaseer):

this should help you solve your problem: http://mathworld.wolfram.com/ParallelVectors.html

OpenStudy (amistre64):

a vector can be thought of as a line in space. lines that are parallel have the same slope. P-Q = vector vector, or slope, from P to Q P(-3,1, 0,5) -Q( 0, 2,1,-3) ------------ V(-3,-1,-1,8) any scaled value of this vector, when attached to any point, will give you a parallel "line". The point given is R: R(-2,1,4,0) + V(-3,-1,-1,8) x=-2-3t y= 1-1t z= 4-1t m= 0+8t Now, we can find any point we want to on this line

OpenStudy (amistre64):

since they want Z to be 8: z= 4-1t = 8 -t = 4 t = -4 will gives us the rest of them